OpenStudy (anonymous):

What is the total time complexity if each instruction takes A ms. The following is a matlab code extract: for i=2:n-3 for j=4:n-1 disp(i*J) end end Show all working if possible thanks.

OpenStudy (anonymous):

The time sued for the outer loop will be something like \[t_{outer} = a \cdot n \cdot t_{inner} + b \] (where a and b are some constants). Since the inner loop also iterates n times as well, \[t_{inner} = c \cdot n \cdot t_{disp} + d\] I assume \[t_{disp} = e \cdot 1\] Now, if we ignore all te constants (like for Big O notation), you'll get something like \[t_{outer} = n \cdot t_{inner} = n \cdot n \cdot t_{disp} = n^2 \cdot 1 = n^2\] So the complexity of the loop is \[O(n^2)\]